A hydrogen stream at 293K and under atmospheric pressure has a dew point of 285K. It is to be further humidified by adding saturated steam at 220kN/m2 through a nozzle at the rate of 2kg steam to 30kg of hydrogen feed. Given that the gas outlet pressure after the mixing process is latm, determine: j) humidity of the hydrogen stream, ii) bumidity of the resultant stream, and iii) temperature of the resultant stream Specific heat of hydrogen-14.6kWkgĶ Specific heat of water (liquid) - 4.18kJ/kg.K Specific heat of water (vapor) -2.0lkJ/kg.i Note: You may need a steam table for the latent heat of water at different pressure and temperature
Answer #1

SHEET. 1 Pad(i) 59.9334 RelaPivo Oim yHo = 0.9862916.SHEET2 Masal oo in Mon al H 1. 972 S833 R mao -o.RS09mm - 26, c4-SHo.RS09% 3. 3355 MagsSHEET 26. 6645 el Pod Ciii)SHEET- 4 335s +[4-84(285-773)+24 뀨·831 4+261+ (793 85) Hw = 8469·6726 54구 k . , 396.4-014 PH Hf 26,6645% 14,6* [T-Q73): 389.3SHEET. S ywF = 5.3355/18 5 335S 13.35?25 Uru on in nof_ О.OR1749 1: ο.091거49adm ,39.19 lo.7A3ST Huc- 19,397.4665% ! 5 + .642104-389 . 301 3486.034 +8449 . 679654++ 54- T293.867414 KDEAR STUDENT ANSWER MAY CHANGE SLIGHTLY DUE TO DIFFERENT CALCULATION METHODS FOR ENTHALPY OF STEAM POSSIBLE. KINDLY CONSIDER IT.

DEAR STUDENT, IF YOU HAVE ANY QUERY KINDLY ASK ME. I SHALL DEFINITELY HELP IN RESOLVING YOUR QUERY. PLEASE ASK ME. THANK YOU.